Class Results Genotype Phenotype Strong Taster Weak Taster Non-taster TT 5 1 Tt 1 10 1 tt 75. (2 pts) What is the frequency of the taster and non-taster allele in the class? Use the most accurate formula.
Answer #1

This shows T dominant allele is responsible for tasting and t is responsible for not tasting.

So, we have to calculate the allelic frequency of T and t allele.

Total people = 25

People with tt phenotype = 7

So, tt genotype frequency = 7/25 = 0.28

For t allele frequency = ✓0.28 = 0.529

Individuals with TT that is homozygous taster = 5+1 = 6

Frequency of TT genotype = 6/25 = 0.24

Therefore, frequency of T allele that is taster allele is ✓0.24 = 0.48

Thus, frequency of taster allele is 0.48

Frequency of non taster allele is 0.52.

Please Hit thumbsup if I could help you. ???